# An LLM-based Recommender System Environment

### v1

@@ A.3.2 Query Template
In the following section, we provide an example prompt and accompanying LLM answer. It is important to note that Listing 4 displays the complete response from the model, not just the rating. During interaction with an RL model, we halt generation after producing the rating.
@@
Listing 4: Example query for rating to the LLM (Vicuna-v1.5-13B). For each user we inject their description, which contains preferences and tastes. Then we provide the movie details: storyline, genres, main actors and vote average.
@@ A.3.3 N-shot Prompting
In this section, we present two illustrative examples (Listing 5 and 6) used for both one-shot and two-shot prompts. These examples adhere to the query template shown in A.3.2, guiding the model on how to formulate responses by considering all features. Additionally, both examples provide comprehensive and detailed answers. One advantage of few-shot prompting is that the examples are given at the start of the prompt. This gives us the chance to use the key-value cache (Pope et al., 2023) without slowing down the environment. Furthermore, these two examples are designed to cover both positive and negative rewards. In the first example, the user rates the queried movie highly, while in the second example, the user rates the queried movie poorly, thereby encompassing a range of user preferences and responses.
@@
Listing 5: First shot example for movies environment.
@@
Listing 6: Second shot example for movies environment.
@@ B.1 Users Generation
We generate the user dataset in the same manner as we did for the users in the movie dataset, by sampling the user features from the same lists using the same method.
@@ B.2 Items
We filtered books from the Amazon Book Dataset (Ni et al., 2019) by removing books that did not have all of the features: categories, description, title, and publication date. We also limit the categories to those with at least 100 books, so we do not get fine-grained categories.
@@ B.3.1 Custom System Prompt
We also experimented with various system prompts, which are predefined text or instructions used to initiate a conversation or request from a user when interacting with a language model. The primary objective was to encourage the model to generate ratings that are less biased and more closely aligned with the information provided to the model. This includes factors such as the user description, the list of books read previously, and the back-cover of the queried book, all of which play a role in shaping the predictions of a model.
@@
Listing 7: An advanced system prompt guiding the model to provide personalized and unbiased movie ratings based on detailed user and movie data.
@@ B.3.2 Query Template
Listing 8: Example query for rating to the LLM. For each user we inject their description, which contains preferences and tastes. Then we provide the movie details: backcover, category, authors and vote average.
@@ B.3.3 N-shot Prompting
In this section, we present two illustrative examples (Listing 9 and 10) used for both one-shot and two-shot prompts. These examples adhere to the query template shown in B.3.2, guiding the model on how to formulate responses by considering all features. Additionally, both examples provide comprehensive and detailed answers. Furthermore, these two examples are designed to cover both positive and negative rewards. In the first example, the user rates the queried book highly, while in the second example, the user rates the queried book poorly, thereby encompassing a range of user preferences and responses.
@@
Listing 9: First shot example for books environment
@@
Listing 10: Second shot example for books environment
Appendix C Extended Ablations Results
Genres/Categories
For the genre test set, we manually created four distinct users for each genre: action, animation, comedy, documentary, family, fantasy, horror, and romance. These users included two women and two men, with one younger individual and one older individual for each gender. In constructing the user descriptions, we ensured that each person consistently rated a specific genre highly (between 8 and 10) while assigning lower ratings to all other genres (between 1 and 5). The asymmetry in high and low ratings is motivated by research done by Ramos et al. (2015), where they analyze rating behavior on IMDB. We then presented these users with a set of 20 movies from their preferred genres and another 20 movies from genres they disliked. Our evaluation metric was the percentage of successful predictions in these scenarios. For the book environment, users are created in the same manner with the exception of their genre preferences, which are specific to book categories rather than movies. The available book categories include fiction, biography, economics, health, philosophy, computer, humor, and drama.
@@ Appendix C Extended Ablations Results
Listing 11: Example description of a user of the genres test set
High/Low Rating
We created eight hand-crafted users: four females and four males, with two young and two elderly individuals of each gender. Within each age group, there is one user who consistently rates items highly and one who consistently rates items low. In this evaluation, we present 160, 20 when using a model based on a paid API (GPT-3.5, GPT-4), items to each of these users, and assess the environment performance by measuring the percentage of successful predictions. The correctness of the environment is determined by its ability to predict high ratings for users whose descriptions explicitly indicate a preference for higher ratings and low ratings for users whose descriptions imply a preference for lower ratings.
@@
Listing 12: Example description of a user of the high/low test set
Collection of Items
To evaluate the movie environment, we took a selection of 22 movie franchises as our test cases. For each franchise, we sample a set of 100 users from our dataset, 50 when using a model based on a paid API (GPT-3.5, GPT-4). To construct the user histories, we included all movies from the respective franchise, except for one, and filled the histories with additional randomly chosen movies. For every user, we designed two distinct queries for the environment. In the first query, all the movies were ones that the user had rated highly in the past. In contrast, in the second query, the user had assigned low ratings to all the movies. Subsequently, we requested a rating for the movie that had been excluded. For the first type of query, we consider the environment to be successful if the user assigns a high rating (consistent with their previous high ratings). In the second type of query, success is determined by the user assigning a low rating (consistent with their previous low ratings). This methodology resulted in the creation of 200 queries for the environment for each franchise. The overall score is calculated based on the percentage of successful predictions across all tests over the 22 different film franchises. We tested the book environment in a similar way, with the only exception that we used 20 book collection.
Similarity to Real Rating Distribution
To calculate the similarity with the true data distribution of MovieLens, we begin by sampling two datasets, $D_{E}$ (for our environment) and $D_{M}$ (for MovieLens), with replacement. Our sampling process is as follows: for the movie environment we randomly select a user and a movie and request the rating. Similarly, for MovieLens, we start by choosing a movie uniformly at random and then choose one of its ratings randomly. Let $D_{E}=\{(m_{1},u_{1},r_{1}),\dots,(m_{N},u_{N},r_{N})\}$ be the dataset sampled from our environment, and let $D_{M}=\{(m_{1}^{\prime},u_{1}^{\prime},r_{1}^{\prime}),\dots,(m_{N^{\prime}}^{\prime},u_{N^{\prime}}^{\prime},r_{N^{\prime}}^{\prime})\}$ be the MovieLens dataset. Where a triplet $(m,u,r)$ represents a user’s rating of movie $m$ with a score of $r$. We compute empirical rating distributions for both MovieLens and the movie dataset from our environment as follows:
@@
$\text{sim}(D_{M},D_{E})=1-\delta(p_{D_{M}}-p_{D_{E}}).$ (4)
For the book environment, we find that existing datasets are too biased toward high ratings to be good candidates, so the aggregated score does not include this test case. In the following section we present additionally results for both the movie and book environments. We show more detailed scores for the different test set of the ablation for the different models, as well as the impact of the different perturbator and retrieval component on the scores of Mistral 7B. Additionally we provide more detail on the impact of different prompting strategies for Vicuna-v1.5-13B in the movie setting
Table 9: Ablation results for the movie environment using the following settings: 0-9 rating scale, 2-shot, custom system prompt, T5-similarity movie retrieval, and perturbation none. For each prompt component we show the aggregated score and specific sub-scores for the various test cases.
@@
Appendix D Details Human Evaluation
@@
Listing 13: Example LLM interaction of the human evaluation
In the human evaluation, we show participants pairs of prompt and LLM response as shown in Listing 13 and ask them to assign a score between 1 and 5 according to the following guideline:
@@
We selected 10 different queries, 4 from the genre test set of the ablation study, 2 from the movie franchise test set of the ablation study, and the remaining 4 we randomly sampled user and movie while manually filling in the previous ratings of the selected users. We successively queried each model with all ten questions using the default configuration (2 shot prompt and custom system prompt with no perturbator), this way we collected 50 different answers, ten for each of the five models. We perform the study by asking 14 participants to score the LLM interactions. In the questionnaire, the order of the questions is randomized for each participants. We then aggregate the score by averaging over each model. Participants in this study were recruited from among our colleagues and provided informed consent to participate. The study was designed with strict adherence to randomized, double-blind procedures to ensure impartiality and reliability of the results. As this user study did not involve tracking or ongoing monitoring of participants, our institutional review board (IRB) determined that formal approval was not required. From the survey (cf. Table 4), we find that users agree most with GPT-4, outperforming all other models. Furthermore, we find that Mistral 7B is the best LLM among open-source models despite only having 7B parameters.
Appendix E RL models
For A2C and PPO we implemented the actor based on the principles of low-rank approximation (Aggarwal et al., 2016). For each user $u$ within the set $U$, we maintain a feature vector $e_{u}$. Similarly, for each movie $m$, we use its feature vector $e_{m}$ and bias $b_{m}$. Additionally, we introduce the movie embedding matrix $E$, and the bias vector $b$. The probability of recommending movie $m$ to user $u$ is calculated as follows:
@@ Appendix E RL models
where $A$ serves as a mask to assign a probability of zero to movies that user $u$ has already viewed. In other words, the entry $A_{m}$ is set to negative infinity if user $u$ has previously watched movie $m$. We employ A2C (Mnih et al., 2016) to train the agent. The actor network, which is responsible for recommending movies, samples actions according to Eq. 5, while the critic consists of a basic two-layer neural network, which takes the user together with the past movie ratings of users as input. We train the model with the default configuration of SB3 (Raffin et al., 2021) for 1.6M steps on SUBER. All parameters are default, except for gamma, which is changed to 0.975. For TRPO and DQN the actor network additionally takes as input the past movie ratings. Also in this case, the models were trained for 1.6M steps with the default configuration of SB3, and all parameters are default except for gamma, which is changed to 0.975.
E.1 Genre Preference
@@
(a) A2C
(b) TRPO
(c) PPO
(d) DQN
Figure 7: User genre preference statistics of top-5 movie recommendations generated by the RL model, trained with different algorithms.
E.2 Recommended Movies Examples
@@ E.3 Environment Performance
In Table 19 we show, for different LLMs, the size and performance (iterations in seconds) of our environment.
Appendix F List of Hobbies and Professions
Please refer to our codebase [^3] https://github.com/giorgiopiatti/SUBER for the following lists:
  • Hobbies: environment/users/datasets/hobby_list.csv
  • Hobbies for childrens: environment/users/datasets
    /hobby_children.csv
  • Professions: environment/users/datasets/job_list.csv

### v2

@@ A.3.2 Query Template
In the following section, we provide an example prompt and accompanying LLM answer. It is important to note that LABEL:listing:example_interaction_1 displays the complete response from the model, not just the rating. During interaction with an RL model, we halt generation after producing the rating.
@@
Listing 6: Example query for rating to the LLM (Vicuna-v1.5-13B). For each user we inject their description, which contains preferences and tastes. Then we provide the movie details: storyline, genres, main actors and vote average.
@@ A.3.3 N-Shot Prompting
In this section, we present two illustrative examples (LABEL:listing:1_shot_prompt_movies and LABEL:listing:2_shot_prompt_movies) used for both one-shot and two-shot prompts. These examples adhere to the query template shown in Section A.3.2, guiding the model on how to formulate responses considering all the features. In addition, both examples provide comprehensive and detailed answers.
One advantage of few-shot prompting is that the examples are given at the start of the prompt. This gives us the opportunity to use the key-value cache [32] without slowing down the environment.
Furthermore, these two examples are designed to cover both positive and negative rewards. In the first example, the user rates the queried movie highly, while in the second example, the user rates the queried movie poorly, thereby encompassing a range of user preferences and responses.
@@
Listing 7: First shot example for movies environment.
@@
Listing 8: Second shot example for movies environment.
@@ B.1 Users Generation
We generate the user dataset in the same way as we did for the users in the movie dataset, by sampling the user features from the same lists using the same method.
@@ B.2 Items
We filter books from the Amazon Book Dataset [28] by removing books that did not have all the features: categories, description, title, and publication date. We also limit the categories to those with at least 100 books, so we do not get fine-grained categories.
@@ B.3.1 Custom System Prompt
We also experimented with various system prompts, which are predefined text or instructions used to initiate a conversation or request from a user when interacting with a language model. The primary objective is to encourage the model to generate ratings that are less biased and more closely aligned with the information provided to the model. This includes factors such as the user description, the list of books read previously, and the back-cover of the queried book, all of which play a role in shaping the predictions of a model.
@@
Listing 9: An advanced system prompt guiding the model to provide personalized and unbiased movie ratings based on detailed user and movie data.
@@ B.3.2 Query Template
Listing 10: Example query for rating to the LLM. For each user we inject their description, which contains preferences and tastes. Then we provide the movie details: backcover, category, authors and vote average.
@@ B.3.3 N-Shot Prompting
In this section, we present two illustrative examples (LABEL:listing:1_shot_prompt_books and LABEL:listing:2_shot_prompt_books) used for both one-shot and two-shot prompts. These examples adhere to the query template shown in Section B.3.2, guiding the model on how to formulate responses by considering all features. Additionally, both examples provide comprehensive and detailed answers.
Furthermore, these two examples are designed to cover both positive and negative rewards. In the first example, the user rates the queried book highly, while in the second example, the user rates the queried book poorly, thereby encompassing a range of user preferences and responses.
@@
Listing 11: First shot example for books environment
@@
Listing 12: Second shot example for books environment
Appendix C Experiment Details
In this appendix we present more details regarding the test cases, showing also the specifics for both implementations of SUBER for movies and books.
Genres/Categories.
For the genre test set, we manually created four distinct users for each genre: action, animation, comedy, documentary, family, fantasy, horror, and romance. These users included two women and two men, with one younger individual and one older individual for each gender. In constructing the user descriptions, we ensured that each person consistently rated a specific genre highly (between 8 and 10) while assigning lower ratings to all other genres (between 1 and 5). The asymmetry in high and low ratings is motivated by research done by Ramos et al. [36], where they analyze rating behavior on IMDB. We then presented these users with a set of 20 movies of their preferred genres and another 20 movies from genres they dislike. Our evaluation metric is the percentage of successful predictions in these scenarios.
For the book environment, users are created in the same manner with the exception of their genre preferences, which are specific to book categories rather than movies. The available book categories include fiction, biography, economics, health, philosophy, computer, humor, and drama.
@@ Appendix C Experiment Details
Listing 13: Example description of a user of the genres test set
High/Low.
We created eight hand-made users: four females and four males, with two young and two elderly individuals of each gender. Within each age group, there is one user who consistently rates items highly and one who consistently rates items low. In this evaluation, we present 160, 20 when using a model based on a paid API (GPT-3.5, GPT-4), items to each of these users, and assess the environment performance by measuring the percentage of successful predictions. The correctness of the environment is determined by its ability to predict high ratings for users whose descriptions explicitly indicate a preference for higher ratings and low ratings for users whose descriptions imply a preference for lower ratings.
@@
Listing 14: Example description of a user of the high/low test set
Collection of Items.
To evaluate the movie environment, we took a selection of 22 movie franchises as our test cases. For each franchise, we sample a set of 100 users from our dataset, 50 when using a model based on a paid API (GPT-3.5, GPT-4). To construct the histories of the users, we included all movies from the respective franchise, except one, and filled the histories with additional randomly chosen movies. For every user, we designed two distinct queries for the environment. In the first query, all the movies were ones that the user had rated highly in the past. In contrast, in the second query, the user had assigned low ratings to all the movies. Subsequently, we requested a rating for the movie that had been excluded. For the first type of query, we consider the environment to be successful if the user assigns a high rating (consistent with their previous high ratings). In the second type of query, success is determined by the user assigning a low rating (in agreement with their previous low ratings).
This methodology resulted in the creation of 200 queries for the environment for each franchise. The overall score is calculated based on the percentage of successful predictions in all tests in the 22 different film franchises.
We tested the book environment in a similar way, with the only exception that we used 20 book collections.
Similarity to Real Rating Distribution.
To calculate the similarity with the true data distribution of MovieLens, we begin by sampling two datasets, $D_{E}$ (for our environment) and $D_{M}$ (for MovieLens), with replacement. Our sampling process is as follows: for the movie environment, we randomly select a user and a movie and request the rating. Similarly, for MovieLens, we start by choosing a movie uniformly at random and then choose one of its ratings randomly.
Let $D_{E}=\{(m_{1},u_{1},r_{1}),\dots,(m_{N},u_{N},r_{N})\}$ be the dataset sampled from our environment, and let $D_{M}=\{(m_{1}^{\prime},u_{1}^{\prime},r_{1}^{\prime}),\dots,(m_{N^{\prime}}^{\prime},u_{N^{\prime}}^{\prime},r_{N^{\prime}}^{\prime})\}$ be the MovieLens dataset. Where a triplet $(m,u,r)$ represents a user’s rating of movie $m$ with a score of $r$.
We compute empirical rating distributions for both MovieLens and the movie dataset from our environment as follows:
@@
$\text{sim}(D_{M},D_{E})=1-\delta(p_{D_{M}},p_{D_{E}}).$ (4)
For the book environment, we find that existing datasets are too biased toward high ratings to be good candidates, so the aggregated score does not include this test case.
Appendix D Extended Ablations Results
In the following section, we present additional results for both the movie and book environments. We show more detailed scores for the different test set of the ablation for the different models, as well as the impact of the different perturbator and retrieval component on the scores of Mistral 7B. Additionally we provide more detail on the impact of different prompting strategies for Vicuna-v1.5-13B in the movie setting
Table 7: Ablation results for the movie environment using the following settings: 0-9 rating scale, 2-shot, custom system prompt, T5-similarity movie retrieval, and perturbation none. For each prompt component, we show the aggregated score and specific subscores for the various test cases.
@@ Appendix D Extended Ablations Results
Appendix E Details Human Evaluation
@@
Listing 15: Example LLM interaction of the human evaluation
In the human evaluation, we show participants pairs of prompt and LLM response as shown in LABEL:listing:user_study_example and ask them to assign a score between 1 and 5 according to the following guideline:
@@
We selected 10 different queries, 4 from the genres test set of the ablation study, 2 from the movie franchise test set of the ablation study, and the remaining 4 we randomly sampled user and movie while manually filling in the previous ratings of the selected users. We successively queried each model with all ten questions using the default configuration (2 shot prompt and custom system prompt with no perturbator), this way we collected 50 different answers, ten for each of the five models. We perform the study by asking 14 participants to score the LLM interactions. In the questionnaire, the order of the questions is randomized for each participants. We then aggregate the score by averaging over each model.
Appendix F RL Models
For A2C and PPO we implemented the actor based on the principles of low-rank approximation [2]. For each user $u$ within the set $U$, we maintain a feature vector $e_{u}$. Similarly, for each movie $m$, we use its feature vector $e_{m}$ and bias $b_{m}$. Additionally, we introduce the movie embedding matrix $E$, and the bias vector $b$. The probability of recommending movie $m$ to user $u$ is calculated as follows:
@@ Appendix F RL Models
where $A$ serves as a mask to assign a probability of zero to movies that user $u$ has already viewed. In other words, the entry $A_{m}$ is set to negative infinity if user $u$ has previously watched movie $m$. We employ A2C [27] to train the agent. The actor network, which is responsible for recommending movies, samples actions according to Equation 5, while the critic consists of a basic two-layer neural network, which takes the user together with the past movie ratings of users as input. We train the model with the default configuration of SB3 [34] for 1.6M steps on SUBER. All parameters are default, except for gamma, which is changed to 0.975. For TRPO and DQN the actor network additionally takes as input the past movie ratings. Also in this case, the models were trained for 1.6M steps with the default configuration of SB3, and all parameters are default except for gamma, which is changed to 0.975.
F.1 Metrics
In this section we outline the different metrics used for the RL model evaluation. In different metrics we need to distinguish if an item is relevant or not, since we don’t have binary interaction, we consider an item relevant for a specific user if the rating given by the user is seven or above. This choice is motivated by Ramos et al. [36].
Mean Average Precision (MAP@10).
@@ F.1 Metrics
For a fixed user, the Average Precision (AP@10) is defined as the average of the precision values calculated at the positions where relevant items appear within the top ten recommendations.
$AP@10=\frac{1}{10}\sum_{k=1}^{10}\text{Precision@}k\cdot R(k),$
where R(k) is 1 if item k is relevant and 0 otherwise. MAP@10 is the average of AP@10 across users.
MRR@$10$ combines precision and relevance across multiple users to provide a comprehensive measure of how well the system ranks relevant items within the top-k recommendations.
Mean Reciprocal Rank (MRR@10).
Mean Reciprocal Rank is used to understand how many recommendations are needed on average before the first relevant item is recommended, and thus measure how quickly a user is satisfied.
For a fixed user, the Reciprocal Rank (RR) is the multiplicative inverse of the rank of the first relevant item. RR@10 is the RR when considering the first ten recommendations (if no relevant item is recommended in the first 10 recommendations, then the value is set to 0). MRR@10 is the average of RR@10 across users.
Personalization (Pers.@10).
To assesses whether a model recommends many of the same items to different users, we use Personalization Pers.@$10$, which is defined as (1 - cosine similarity) between user’s lists of top 10 recommendations.
Let $M$ be the user-item interaction matrix, where $M_{u,i}=1$ if user $u$ has been recommended item $i$ in the first 10 recommendations and $M_{u,i}=0$ otherwise.
[…]
Let $\mathcal{P}$ be the set of all user pairs $(u,v)$ such that $u>v$.
The personalization metric is defined as:
[…]
A high personalization score indicates that recommendations vary significantly among different users, suggesting that the model offers a more personalized experience to each user.
Liked Genres.
Each user in the training dataset has both preferred and disliked movie genres. The trained RL model generates a list of the top 5 movie recommendations for each user. The recommendations are classified into three categories: liked (movies matching preferred genres and excluding disliked ones), disliked (movies with disliked genres and no preferred ones), and neutral (remaining recommendations). “Liked Genres” is the percentage of recommendations that fall into the first category. For each user $u$ we define $L(u)$ to be the percentage of movies that fall into the liked genre category in the top-5 recommendation, then the liked genre metric is the average over the dataset of users:
Linked $\text{Linked Genres}=\frac{1}{|U|}\sum_{u\in U}L(u).$
This metric measures how well the recommender system makes genre-specific recommendations.
F.2 Genre Preference
In this section, we present the user-genre preference statistics for the top-5 recommendations generated by A2C, TRPO, PPO, and DQN models. The results show that, for all models, the majority of recommendations fall into the liked category. By comparing Figure 4 with Figures 8, 9, 10 and 11, we observe a positive correlation between model rewards and the percentage of top-5 recommended movies in the "liked" category. In other words, the better the model is at recommending movies that align with users’ preferred genres, the higher the reward it achieves in the environment.
Moreover, as shown in Figures 8, 9, 10 and 11 the RL recommender model is able to learn the dynamics of genre preferences of users, mainly recommending movies that fall into the favored genres of users. It should be noted that recommending a neutral movie can be a valid strategy, especially if it is a highly praised or outstanding movie.
Figure 8: User genre preference statistic of top-5 movie recommendations generated by the RL model, trained with A2C.
Figure 9: User genre preference statistic of top-5 movie recommendations generated by the RL model, trained with TRPO.
Figure 10: User genre preference statistic of top-5 movie recommendations generated by the RL model, trained with PPO.
Figure 11: User genre preference statistic of top-5 movie recommendations generated by the RL model, trained with DQN.
F.3 Recommended Movies Examples
For a set of random users, we interact with the trained RL model, in Tables 14, 15 and 16 with embedding dim 32, and show the first 5 recommended movies.
@@ F.3 Recommended Movies Examples
Table 15: At the top we show the personal interest of Ava. At the bottom, we show the title and genres of the first 5 recommended movies.
@@
Table 16: At the top we show the personal interest of Maya. At the bottom, we show the title and genres of the first 5 recommended movies.
@@ F.4 Environment Performance
Appendix G List of Hobbies and Professions
Table 18: List of children’s hobbies
Hobby-name
Drawing and painting
Playing piano
Playing guitar
Playing violin
Playing flute
Playing drums
Dancing
Reading books
Writing stories
Board games
Card games
Gardening
Cooking
Backing
Building with Lego
Collecting stamps
Collecting coins
Collecting cards
Photography
Learning magic tricks
Soccer
Basketball
Swimming
Volleyball
Tennis
Acting
Singing
Puppetry
Birdwatching or nature exploration
Science experiments
Playing video games
Origami
Learning a new language
Table 19: Jobs list
Jobs
Account Manager
Accountant
Actor
Actuary
Administrator
Advertising Executive
Aerospace Engineer
Aerospace Technician
Air Traffic Controller
Animal Trainer
Architect
Archivist
Art Director
Artist
Auctioneer
Auto Mechanic
Baggage Handler
Bailiff
Baker
Banker
Barber
Barber Shop Owner
Barista
Bartender
Benefits Administrator
Bicycle Mechanic
Biologist
Blacksmith
Boat Captain
Bodyguard
Bookkeeper
Botanical Illustrator
Botanist
Brewery Worker
Bricklayer
Broadcast Technician
Building Inspector
Bus Driver
Bus Mechanic
Butcher
CIO (Chief Information Officer)
Cabin Crew
Cake Decorator
Call Center Operator
Car Salesperson
Carpenter
Cartographer
Cashier
Casino Dealer
Caterer
Chaplain
Chauffeur
Chef
Chemical Engineer
Chemist
Chief Financial Officer (CFO)
Chimney Sweep
Chiropractor
Civil Engineer
Claims Adjuster
Cleaner
Clown
Coach
Coachbuilder
Commercial Pilot
Composer
Computer Programmer
Computer Systems Analyst
Concierge
Conservationist
Table 20: Jobs list
Construction Worker
Cost Estimator
Counselor
Courier
Court Reporter
Craftsperson
Cruise Ship Captain
Cryptographer
Curator
Customer Service Representative
Dairy Farmer
Dancer
Data Analyst
Data Entry Operator
Database Administrator
Demolition Worker
Dental Hygienist
Dentist
Designer
Desktop Publisher
Detective
Detective Inspector
Dialysis Technician
Diesel Mechanic
Dietician
Digital Marketer
Dispatch Operator
Doctor
Dog Trainer
Door-to-Door Salesperson
Dressmaker
Drummer
Dry Cleaner
Economist
Economist
Electrician
Engineer
Event Planner
Farmer
Fashion Designer
Firefighter
Flight Attendant
Florist
Forensic Scientist
Gardener
Geologist
Graphic Designer
Hairdresser
Historian
Hotel Manager
Human Resources Manager
Illustrator
Industrial Designer
Insurance Agent
Interior Designer
Interpreter
Janitor
Journalist
Judge
Laboratory Technician
Lawyer
Librarian
Lifeguard
Linguist
Makeup Artist
Table 21: Jobs list
Locksmith
Manager
Marketing Specialist
Massage Therapist
Mechanic
Medical Assistant
Meteorologist
Model
Musician
Nanny
Nurse
Nutritionist
Occupational Therapist
Optician
Painter
Paramedic
Pharmacist
Photographer
Physical Therapist
Physician Assistant
Pilot
Plumber
Police Officer
Politician
Postal Worker
Producer
Professor
Psychologist
Public Relations Specialist
Real Estate Agent
Receptionist
Reporter
Research Scientist
Sales Representative
Scientist
Security Guard
Singer
Social Media Manager
Social Worker
Software Developer
Sound Engineer
Speech Therapist
Sports Coach
Statistician
Stockbroker
Surveyor
Tailor
Teacher
Technical Writer
Technician
Therapist
Tour Guide
Translator
Travel Agent
Truck Driver
UI/UX Designer
Veterinarian
Video Editor
Waiter/Waitress
Web Developer
Welder
Writer
Yoga Instructor
Zookeeper